40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 10 01:40:40 2011, Koji, Update, SUS, ETMY exploration 
    Reply  Wed Aug 10 02:29:38 2011, Jenne, Update, SUS, ETMY exploration 
Message ID: 5163     Entry time: Wed Aug 10 01:40:40 2011     Reply to this: 5164
Author: Koji 
Type: Update 
Category: SUS 
Subject: ETMY exploration 

[Rana Koji Jenne Jamie]

- The situation of the ETMY suspension is improved.
- The damping servos except for Pitch are now functional.
- We intentionally turned off the damping servos for the matrix measurements.


- Opened the light door of the ETMY chamber.

- We set up the CDS SUS lockin:

        Excite UL/UR/LL/LR equally [by setting the output matrix (1, 1, 1, 1, 0)] at 3.12Hz with 2000 cnts
        Put the OSEM PD outputs into lockin one by one. For the image rejection, 0.1Hz 4th order LPF has been used though we like to use a faster settling LPF.

- Found only UL was responding to the excitation. After fitzing with the cables and connectors, it was found that the DAC card was loose from the bus.
  By pushing the card the responses have been back. [Note we had the reboot of c1iscey almost at the same time.]

- Checked the response in the I channel of the lockin.
        UL -8ish / UR +7ish / LR +5ish / LL +2ish

- Tweaked LL sensor to get better response ==> in vain. Decided to move the lower OSEM plate for the better positioning of the LR/LL.
- Got reasonable (+5ish) response for LL.

- Confirmed that the POS/YAW/SIDE damping works with positive gains. PITCH did not work with the negative gain (but that could be a good sign.)

- Let the suspension freely swinging for a while (~30min). Checked the side/pos separation. They are not perfect but seemed diagonalizable.

- Closed the light door.

- Jenne will make a better kick/free-swing test later.

ELOG V3.1.3-